Solve the system. Thank you for your help.
6x+5y=2
10x-7y=11

Answer by stanbon(75887)

6x+5y=2
10x-7y=11
--------------
Modify for elimination of "x":
60x + 50y = 20
60x - 42y = 66
-----
Subtract and solve for "y":
92y = -46
y = -1/2
---------------------
Modify for elimination of "y":
42x + 35y = 14
50x - 35y = 55
----------------
92x = 69
x = 3/4
================
Cheers,
Stan H.
---------------